OpenStudy (anonymous):

a telephone company charges $1.80 for a 12 minute long distance call, and $2.80 for a 20 minute call to the same number on the same day and time. find a linear equation that expresses the total charges c in terms of t...... last question that i can't figure out!! @raffle_snaffle

OpenStudy (tboi):

Ok first, what is the first step you have to do?

OpenStudy (faiqraees):

Total charges = Charges for 12 minute+ charges for 20 minute 1.8 +2.8 = 1.8 +2.8 4.6 = 12t +20t 4.6=32t Is this what you're asking?